# Normal approximations for the multivariate inverse Gaussian distribution and asymmetric kernel smoothing on $d$-dimensional half-spaces

### v4

@@ 4 Exact simulation algorithm
Compared with Minami’s algorithm (Minami, 2003, Theorem 2), which simulates the first-hitting location on a hyperplane of a $d$-variate Brownian motion with correlated components, the procedure described below Proposition 4.1 offers two main advantages: it delivers exact samples and achieves significantly higher speed. Minami’s approach is only approximate because it relies on discretizing the Brownian paths, leading to slow computing times even in low dimension. In contrast, the exact algorithm involves inverting the covariance matrix $\mathbf{\Omega}$ and performing cross-products in quadratic forms, which yields a computational complexity of $\mathcal{O}(d^{3}+n\,d)$, linear in the sample size $n$. Table 1 reports the median computing time for generating $n=100$ random vectors from a MIG distribution with randomized parameters, across different dimensions $d$. The Brownian motion simulations used a time discretization step of $10^{-3}$ to approximate first-hitting locations. The table demonstrates that the proposed method is several orders of magnitude faster than Minami’s.
Table 1: Median time (out of 25 replications) in milliseconds for the generation of $n=100$ observations from a MIG distribution with randomized parameters as a function of $d$.
Dimension ($d$) 2 3 5 10
Exact 0.52 0.83 0.93 3.10
Minami 3,060 3,823 5,664 14,904
@@ Appendix B Proofs of the asymptotics for the MIG kernel density estimator
Therefore, by fixing $\varepsilon\in(0,\infty)$, taking the expectation on both sides of (B.2), and then using (B.1), one can bound the bias as follows:

### v5

@@ 4 Exact simulation algorithm
Compared with Minami’s algorithm (Minami, 2003, Theorem 2), which simulates the first-hitting location on a hyperplane of a $d$-variate Brownian motion with correlated components, the procedure described below Proposition 4.1 offers two main advantages: it delivers exact samples and achieves significantly higher speed. Minami’s approach is approximate because it relies on discretizing the Brownian paths, leading to slow computing times even in low dimension. In contrast, the exact algorithm involves inverting the covariance matrix $\mathbf{\Omega}$ and performing cross-products in quadratic forms, which yields a computational complexity of $\mathcal{O}(d^{3}+n\,d^{2})$, linear in the sample size $n$.
Table 1 reports the mean computation time and corresponding standard deviation for generating $n=100$ MIG random vectors with randomized parameters, across different dimensions $d$. Minami’s Brownian-motion method was run with a time-discretization step of $10^{-3}$ to approximate first-hitting locations.
Table 1: Mean time (standard deviation) in milliseconds over 25 replications for generating $n=100$ observations from a MIG distribution with randomized parameters, as a function of $d$.
Method $d=2$ $d=4$ $d=8$ $d=16$ $d=32$
Exact 0.80 (0.27) 1.31 (0.37) 1.61 (0.36) 1.98 (0.42) 3.22 (0.40)
Minami 259 (45) 2275 (984) 14228 (5024) 38373 (21001) 197823 (366493)
The table shows that the exact generator is up to five orders of magnitude faster than Minami’s Brownian-motion algorithm; the same pattern is visible in Figure 10, which plots the mean times with 95% confidence intervals and log-log least-squares lines.
Figure 10: Computation time for rmig generation versus dimension. Points show the mean run time across 25 repetitions, and vertical bars 95% Wald-based confidence intervals for the mean. Dashed lines give ordinary least-squares fits in log-log scale. The exact algorithm is compared with Minami’s Brownian motion method.
@@ Appendix B Proofs of the asymptotics for the MIG kernel density estimator
Therefore, by fixing $\varepsilon\in(0,\infty)$, taking the expectation on both sides of (B.2), and then using (B.1), one can bound the recentered bias as follows: